### v4

# Physics-Constrained Learning of Energy-Preserving Stencils for Maxwell’s Equations

Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Starting from high-fidelity spectral training data, we learn a linear convolution stencil that approximates the spatial derivative in Maxwell’s equations. The stencil coefficients are obtained from a convex quadratic program with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol. We show that the learned operator has a purely imaginary symbol, induces purely oscillatory modal dynamics, and preserves the discrete Maxwell energy exactly under Crank–Nicolson time-stepping. Numerical experiments with smooth traveling-wave solutions demonstrate that the energy-constrained, ADMM-learned stencil with radius $R=1$ achieves approximately second-order spatial convergence in the PDE error, matching the expected order of the standard centered finite-difference stencil. We compare projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point solver for the resulting quadratic program, and we illustrate extensions to nonstandard target operators and noisy training data. Overall, the learned stencils achieve accuracy comparable to classical central differences while exactly preserving a discrete electromagnetic energy, making them viable drop-in, structure-preserving spatial operators for finite-difference time-domain methods and data-driven PDE solvers.
1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [25, 26]. Classical schemes, such as the Yee method [28], are derived directly from the continuous equations: spatial derivatives are approximated using fixed central-difference stencils, and stability is enforced via analytical CFL conditions. These schemes are carefully designed so that the discrete curl operator is skew-adjoint, which yields exact semi-discrete energy conservation and compatible discrete divergence constraints.
Recently, there has been growing interest in data-driven discretizations, in which discrete operators are learned from numerical or experimental data, for example via regression or neural networks. Examples include data-driven finite-difference discretizations for Burgers’ equation and the Navier–Stokes equations [1], data-driven closures for turbulent flows [21, 14], and sparse-regression or neural approaches for discovering governing equations and hidden dynamics [24, 22, 7]. These methods typically aim to minimize approximation error on a chosen training distribution, but they do not, in general, explicitly enforce structural properties such as stability, conservation, or energy balance.
For Maxwell’s equations, this is a significant limitation. Classical FDTD schemes are constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level, ensuring that numerical solutions respect key physical invariants. Learned operators that ignore these structures risk exhibiting uncontrolled numerical dissipation or growth, even if they fit the training data well.
In this work we address this gap in a simple but concrete setting: the one-dimensional Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
$D:\mathbb{R}^{N}\to\mathbb{R}^{N},$
represented as a convolution stencil, directly from data. We impose linear constraints on the stencil that enforce skew-adjointness of $D$ with respect to the discrete $L^{2}$ inner product. This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, independently of the details of the training data or the optimization procedure.
Our approach is inspired by two lines of work. On the one hand, classical structure-preserving discretizations for Maxwell and related PDEs use staggered grids and skew-adjoint discrete curl operators to obtain discrete energy identities and long-time stability [4, 8, 19]. High-order and operator-splitting variants maintain these properties by design [4, 3]. On the other hand, data-driven PDE schemes learn local stencils or coarse-grid operators from high-fidelity simulations, often achieving accuracy comparable to or better than hand-crafted schemes on specific tasks [1, 11, 6]. However, these learned operators usually encode physical structure only implicitly, through the choice of loss function, training data, or architecture, rather than via hard constraints.
Related ideas also appear in recent neural-operator surrogates for electromagnetics and other time-dependent PDEs. Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material-dispersion structure [17, 13]. Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [18, 10]. Other frameworks embed conservation laws via soft constraints or corrective layers [16, 15, 27]. These approaches can significantly improve stability and physical fidelity, but they generally enforce conservation or energy balance only approximately, depending on optimization and hyperparameters.
Building on this context, our main contributions are as follows. We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data. The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, subject to linear constraints that encode skew-adjointness of the stencil coefficients together with simple box bounds. We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, which in turn yields an exact semi-discrete electromagnetic energy identity and expresses the CFL condition directly in terms of the stencil’s Fourier symbol $\mu(\theta)$. This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law. We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator via this symbol. We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations. We also illustrate how the same constrained learning framework can recover nonstandard, coarse-grid effective operators and regularize synthetic spectral training data contaminated by noise or mild inconsistencies. In particular, the ADMM-learned radius-$R=1$ stencil exhibits approximately second-order spatial convergence for smooth traveling-wave solutions, matching the expected order of the standard centered finite-difference stencil it approximates. The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and the interior-point solver produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime. Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation. To the best of our knowledge, this is the first data-driven convolution stencil for the one-dimensional Maxwell system whose learned coefficients are constrained to guarantee a semi-discrete electromagnetic energy identity.
The paper is organized as follows. Sections 2–3 introduce the semi-discrete 1D Maxwell system, derive the corresponding discrete energy identity, and analyze the numerical wave speed and CFL restriction associated with the learned operator. Sections 4–6 describe the learning formulation and optimization algorithms, and present numerical experiments that evaluate accuracy, energy behavior, dispersion, and robustness to nonstandard targets and noisy training data.
2 Semi-discrete Maxwell system and energy identity
We consider the 1D Maxwell system in vacuum, written in non-dimensional form with $\varepsilon_{0}=\mu_{0}=1$:
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E,$ (1)
on the periodic domain $[0,L]$.
We discretize space using a uniform grid $x_{i}=i\Delta x$, $i=0,\dots,N-1$, with $\Delta x=L/N$, and collect the nodal values of the electric and magnetic fields into vectors
$\mathbf{E}(t)=(E_{0}(t),\dots,E_{N-1}(t))^{\top},\qquad\mathbf{H}(t)=(H_{0}(t),\dots,H_{N-1}(t))^{\top}.$
Let $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$ denote a linear discrete derivative operator. In this work, $D$ is a convolution operator defined by a stencil $w\in\mathbb{R}^{2R+1}$,
$(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell},\qquad i=0,\dots,N-1,$ (2)
with periodic wrap-around in the index $i+\ell$. The semi-discrete Maxwell system then reads
$\frac{d}{dt}\mathbf{E}(t)=D\mathbf{H}(t),\qquad\frac{d}{dt}\mathbf{H}(t)=D\mathbf{E}(t).$ (3)
We equip $\mathbb{R}^{N}$ with the standard discrete $L^{2}$ inner product
$\langle\mathbf{u},\mathbf{v}\rangle:=\Delta x\sum_{i=0}^{N-1}u_{i}v_{i},\qquad\|\mathbf{u}\|^{2}:=\langle\mathbf{u},\mathbf{u}\rangle.$ (4)
The corresponding discrete electromagnetic energy is
$\mathcal{E}(t):=\frac{1}{2}\|\mathbf{E}(t)\|^{2}+\frac{1}{2}\|\mathbf{H}(t)\|^{2}.$ (5)
Theorem 1 (Semi-discrete energy identity).
Let $D$ be any linear operator on $\mathbb{R}^{N}$, and let $\mathbf{E}(t),\mathbf{H}(t)$ satisfy (3). Then
$\frac{d}{dt}\mathcal{E}(t)=\langle D\mathbf{H}(t),\mathbf{E}(t)\rangle+\langle D\mathbf{E}(t),\mathbf{H}(t)\rangle.$ (6)
In particular, if $D$ is skew-adjoint with respect to (4), i.e.,
$\langle D\mathbf{u},\mathbf{v}\rangle=-\langle\mathbf{u},D\mathbf{v}\rangle\quad\forall\,\mathbf{u},\mathbf{v},$ (7)
then the discrete energy is exactly conserved:
$\frac{d}{dt}\mathcal{E}(t)=0.$ (8)
Proof.
Differentiating (5) in time and using (3) gives
$\frac{d}{dt}\mathcal{E}(t)$ $=\left\langle\frac{d}{dt}\mathbf{E}(t),\mathbf{E}(t)\right\rangle+\left\langle\frac{d}{dt}\mathbf{H}(t),\mathbf{H}(t)\right\rangle$
$=\langle D\mathbf{H}(t),\mathbf{E}(t)\rangle+\langle D\mathbf{E}(t),\mathbf{H}(t)\rangle,$
which is (6). If $D$ is skew-adjoint, then
$\langle D\mathbf{H},\mathbf{E}\rangle=-\langle\mathbf{H},D\mathbf{E}\rangle=-\langle D\mathbf{E},\mathbf{H}\rangle,$
so the two terms in (6) cancel and $\frac{d}{dt}\mathcal{E}(t)=0$. ∎
2.1 Skew-symmetric convolution operators and stencil constraints
We now specialize to convolution operators of the form (2) with periodic boundary conditions. Let $w=(w_{-R},\dots,w_{-1},w_{0},w_{1},\dots,w_{R})^{\top}$ be the stencil.
Lemma 1.
Let $D$ be defined by (2) with periodic wrap-around. Then the matrix entries of $D$ satisfy
$D_{ij}=w_{i-j},\qquad i,j=0,\dots,N-1,$ (9)
where the difference iji-j is taken modulo $N$.
Proof.
By definition,
$(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell}=\sum_{j=0}^{N-1}w_{i-j}\,u_{j},$
where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$; see, e.g., [4]. ∎
Theorem 2 (Characterization of skew-symmetric stencils).
Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
$w_{0}=0,\qquad w_{-\ell}=-w_{+\ell},\quad\ell=1,\dots,R.$ (10)
Proof.
By Lemma 1 the entries of $D$ are $D_{ij}=w_{i-j}$, so
$(D^{\top})_{ij}=D_{ji}=w_{j-i}.$
The condition $D^{\top}=-D$ is therefore equivalent to
[…]
or, writing $\ell=i-j$,
w=w+for all $_{-\ell}=-w_{+\ell}\quad\text${for all $\ell$.
In particular, taking $\ell=0$ yields $w_{0}=-w_{0}$, hence $w_{0}=0$, and for $\ell=1,\dots,R$ we obtain $w_{-\ell}=-w_{+\ell}$. Conversely, if (10) holds, then the above computation shows $D^{\top}=-D$, i.e., $D$ is skew-adjoint. ∎
We encode the conditions (10) as a set of linear constraints on the stencil vector $w$. Let $C\in\mathbb{R}^{(R+1)\times(2R+1)}$ be defined by
$(Cw)_{0}=w_{0},\qquad(Cw)_{\ell}=w_{-\ell}+w_{+\ell},\quad\ell=1,\dots,R,$
and let $d\in\mathbb{R}^{R+1}$ be the zero vector. Then the constraint $Cw=d$ is simply
$w_{0}=0,\qquad w_{-\ell}+w_{+\ell}=0,\quad\ell=1,\dots,R,$ (11)
which is equivalent to skew-symmetry.
Corollary 1.
Let $C$ and $d$ be defined in (11). Then $Cw=d$ if and only if the stencil $w$ satisfies (10), i.e., if and only if the corresponding convolution operator $D$ is skew-adjoint.
Proof.
Immediate from (11) and Theorem 2. ∎
Corollary 1 shows that the linear constraint $Cw=d$ used in our optimization problem is exactly the condition required for semi-discrete energy conservation via Theorem 1.
3 Wave speed and CFL for the learned operator
The discrete curl operator $D$ constructed in Section 2.1 is a periodic convolution with stencil $w=(w_{-R},\dots,w_{0},\dots,w_{R})$ satisfying the skew-symmetry constraints of Theorem 2. For such an operator it is convenient to work with its Fourier symbol
$\mu(\theta):=\sum_{k=-R}^{R}w_{k}\,\mathrm{e}^{\mathrm{i}k\theta},\qquad\theta\in[-\pi,\pi],$ (12)
which characterizes the response of $D$ to discrete Fourier modes $\mathrm{e}^{\mathrm{i}j\theta}$.
For the semi-discrete Maxwell system
$\frac{d}{dt}\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix}=A\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix},\qquad A=\begin{bmatrix}0&D\\ D^{T}&0\end{bmatrix}$,
the eigenvalues of $A$ are purely imaginary and can be written as $\lambda(\theta)=\pm\mathrm{i}\,|\mu(\theta)|$ for $\theta\in[-\pi,\pi]$. In particular, the quantity $|\mu(\theta)|$ plays the role of a numerical wave speed for the mode $\theta$, and the maximal value
$c_{\max}:=\max_{\theta\in[-\pi,\pi]}|\mu(\theta)|$ (13)
is the largest speed supported by the learned operator. For the standard centered difference stencil of width three one recovers $c_{\max}\sim h^{-1}$, corresponding to the usual CFL restriction for explicit wave propagation schemes.
If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [23, 20, 12]. Substituting $\omega=\pm|\mu(\theta)|$ and maximizing over $\theta$ yields the CFL condition
$\Delta t\;\leq\;\frac{2}{c_{\max}},$ (14)
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator. In the present work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator. A systematic comparison with leapfrog using the derived CFL bound is left for future work.
3.1 Crank–Nicolson time integration
For the actual time-dependent simulations in Section 6 we use the Crank–Nicolson (CN) method applied to the semi-discrete system
$\frac{d\mathbf{U}}{dt}=A\mathbf{U},\qquad\mathbf{U}=\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix},\qquad A^{T}=-A.$
The CN update reads
$\frac{\mathbf{U}^{n+1}-\mathbf{U}^{n}}{\Delta t}=A\,\frac{\mathbf{U}^{n+1}+\mathbf{U}^{n}}{2},$ (15)
which coincides with the implicit midpoint rule for linear systems. Since $A$ is skew-adjoint, the discrete energy
$\mathcal{E}^{n}:=\tfrac{1}{2}\|\mathbf{U}^{n}\|_{2}^{2}=\tfrac{1}{2}\bigl(\|\mathbf{H}^{n}\|_{2}^{2}+\|\mathbf{E}^{n}\|_{2}^{2}\bigr)$
is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [8, 9] for background on such structure-preserving schemes.
In particular, CN does not require the CFL restriction (14) for stability or energy conservation in this linear, skew-adjoint setting. In the numerical experiments we nevertheless choose a time step $\Delta t$ proportional to the grid spacing $h$ and small enough that (14) would be satisfied for the corresponding centered-difference operator. This ensures that temporal discretization errors remain small relative to the spatial error while allowing us to interpret the learned stencil in terms of an effective wave speed $c_{\max}$.
4 Data-driven stencil learning
We now formulate the quadratic optimization problem used to learn stencils from data generated by a high-fidelity discretization.
4.1 Training data from spectral 1D Maxwell
We generate training data by sampling random Maxwell states and computing their time derivatives exactly (up to spectral discretization error). We work on the periodic domain $[0,L]$ with $N$ grid points and spacing $\Delta x=L/N$. For each training sample we draw a random truncated Fourier series
$E(x)$ $=\sum_{m=1}^{m_{\max}}a_{m}\sin\left(\frac{2\pi mx}{L}+\phi_{m}\right),$
$H(x)$ $=\sum_{m=1}^{m_{\max}}b_{m}\sin\left(\frac{2\pi mx}{L}+\psi_{m}\right),$
with amplitudes $a_{m},b_{m}$ and phases $\phi_{m},\psi_{m}$ drawn from simple Gaussian and uniform distributions. We then compute the spatial derivatives $\partial_{x}E$ and $\partial_{x}H$ spectrally using FFTs:
[…]
where $k$ is the discrete wavenumber. Finally, we obtain the time derivatives from the PDE (1):
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E.$
We denote by
$U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
the array of training states, where $U[s,0,i]=E^{(s)}(x_{i})$ and $U[s,1,i]=H^{(s)}(x_{i})$ denote the electric and magnetic fields of the $s$-th training sample evaluated at the grid point $x_{i}$. Similarly,
$\frac{dU}{dt}\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
stores the corresponding time derivatives, with $(dU/dt)[s,0,i]=\partial_{t}E^{(s)}(x_{i})$ and $(dU/dt)[s,1,i]=\partial_{t}H^{(s)}(x_{i})$. Here $n_{\mathrm{sims}}$ denotes the number of independent training samples (or snapshots), each consisting of a randomly generated Maxwell state $(E,H)$ and its corresponding time derivative. The spatial domain $[0,L]$ is discretized using $N$ uniformly spaced grid points $x_{i}=i\Delta x$, $i=0,\dots,N-1$, where $\Delta x=L/N$.
4.2 Least-squares formulation
Given training pairs $(\mathbf{E},\mathbf{H})$ and $(\partial_{t}\mathbf{E},\partial_{t}\mathbf{H})$, we wish to learn a single convolution stencil $w$ such that
$D\mathbf{H}\approx\partial_{x}\mathbf{H}=\partial_{t}\mathbf{E},\qquad D\mathbf{E}\approx\partial_{x}\mathbf{E}=\partial_{t}\mathbf{H},$
in a least-squares sense. Because we assume periodic boundary conditions, the stencil is applied identically at every grid point, with indices interpreted modulo $N$.
For each grid index $i$ we extract local patches
$\mathbf{h}_{i}=(H_{i-R},\dots,H_{i+R})^{\top},\qquad\mathbf{e}_{i}=(E_{i-R},\dots,E_{i+R})^{\top},$
and define scalar targets
$y^{(H)}_{i}=\partial_{t}E_{i},\qquad y^{(E)}_{i}=\partial_{t}H_{i}.$
The desired relationships are
$w^{\top}\mathbf{h}_{i}\approx y^{(H)}_{i},\qquad w^{\top}\mathbf{e}_{i}\approx y^{(E)}_{i}$.
Stacking all patches and targets over space and all simulations, we obtain a linear system
$Aw\approx b,$ (16)
where each row of $A$ is either $\mathbf{h}_{i}^{\top}$ or $\mathbf{e}_{i}^{\top}$, and $b$ collects the corresponding targets. This leads to the unconstrained least-squares problem
$\min_{w\in\mathbb{R}^{2R+1}}\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (17)
where $\lambda\geq 0$ is a small Tikhonov regularization parameter.
4.3 Energy constraint and box bounds
To enforce skew-adjointness and hence discrete energy conservation, we impose the linear constraint $Cw=d$ with $C$ and $d$ defined as in Section 2.1. By Corollary 1, any stencil satisfying $Cw=d$ defines a skew-adjoint convolution operator $D$ and therefore an energy-conserving semi-discrete Maxwell system.
We also impose simple box bounds $|w_{\ell}|\leq M$, i.e.,
$M\leq w_{\ell}\leq M,\qquad\ell=-R,\dots,R,$ (18)
where $M>0$ is chosen large enough to contain the true derivative stencil but small enough to control numerical conditioning.
Our final optimization problem is the convex quadratic program
$\min_{w\in\mathbb{R}^{2R+1}}$ $\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (19)
s.t. $Cw=d,$
$-M\leq w\leq M.$
This problem is strictly convex for $\lambda>0$ and admits a unique solution. For our first-order methods we use a slightly simplified variant of (19) in which only the linear energy constraint is enforced. Specifically, projected gradient (PG) and Nesterov-accelerated gradient (NAG) are applied to the equality-constrained problem
$\min_{w\in\mathbb{R}^{2R+1}}$ $\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (20)
s.t. $Cw=d,$
while the full equality-plus-box-constrained problem (19) is handled by ADMM and the interior-point solver in CVXPY. In our experiments the solutions of (20) produced by PG and NAG remain well inside the box $M,M]^{2R+1}$, so the difference between (20) and (19) is negligible in practice. In the presence of noisy or mildly inconsistent training data, the least-squares system can become ill-conditioned and produce stencils with very large coefficients; the Tikhonov term and box bounds mitigate this behavior, as illustrated in Section 6.7.
5 Convex solvers
We briefly summarize the algorithms used to solve (19). All methods operate on the same design matrix $A$, constraint matrix $C$, and bounds.
5.1 Projected gradient methods
Our first two solvers are projected gradient (PG) and Nesterov-accelerated gradient (NAG), applied to the equality-constrained problem (20). In particular, for PG/NAG we enforce only the linear energy constraint $Cw=d$ and do not explicitly impose the box bounds $M\leq w\leq M$. In our experiments the PG/NAG iterates remain strictly inside the box, so the resulting solutions agree (to numerical tolerance) with those of the fully constrained problem (19).
Define the feasible affine subspace
$\mathcal{C}_{\mathrm{eq}}=\{w\in\mathbb{R}^{2R+1}:Cw=d\}.$
The orthogonal projection onto $\mathcal{C}_{\mathrm{eq}}$ with respect to the Euclidean inner product is given explicitly by
$P_{\mathcal{C}_{\mathrm{eq}}}(z):=z-C^{\top}(CC^{\top})^{-1}(Cz-d),\qquad z\in\mathbb{R}^{2R+1},$ (21)
assuming that $C$ has full row rank.
Let
$f(w):=\tfrac{1}{2}\|Aw-b\|_{2}^{2}+\tfrac{\lambda}{2}\|w\|_{2}^{2},\qquad\nabla f(w)=A^{\top}(Aw-b)+\lambda w.$
We choose a stepsize $\alpha=1/L$, where one may take $L=\|A\|_{2}^{2}+\lambda=\|A^{\top}A\|_{2}+\lambda$.
For PG, the update is
$w^{k+1}=P_{\mathcal{C}_{\mathrm{eq}}}\!\left(w^{k}-\alpha\nabla f(w^{k})\right),$
with $P_{\mathcal{C}_{\mathrm{eq}}}$ given by (21).
For NAG we use the extrapolated point
$y^{k}=w^{k}+\beta_{k}(w^{k}-w^{k-1}),$
and then update
$w^{k+1}=P_{\mathcal{C}_{\mathrm{eq}}}\!\left(y^{k}-\alpha\nabla f(y^{k})\right),$
with the standard parameters
$t_{k+1}=\tfrac{1}{2}\!\left(1+\sqrt{1+4t_{k}^{2}}\right),\qquad\beta_{k}=\frac{t_{k}-1}{t_{k+1}},\qquad t_{0}=1.$
We track the objective $f(w^{k})$, wall-clock time, and the equality residual $\|Cw^{k}-d\|_{2}$ at each iteration.
Remark 1.
Projecting sequentially onto $\{Cw=d\}$ and the box $M,M]^{2R+1}$ is not, in general, the exact Euclidean projection onto their intersection. If the box constraint becomes active and one wishes to enforce both constraints within a projected-gradient framework, one should compute $P_{\{Cw=d\}\cap[-M,M]^{2R+1}}$ using an intersection projection method such as Dykstra’s algorithm, or solve the projection step as a small quadratic program.
5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [5] scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
$w\in\mathbb{R}^{2R+1},\qquad z\in[-M,M]^{2R+1},\qquad w=z,\qquad Cw=d.$
The augmented Lagrangian splits into a quadratic $w$-problem with equality constraints and a simple proximal $z$-update given by box projection. The $w$-update reduces to solving a KKT system of the form
$\begin{bmatrix}A^{\top}A+(\lambda+\rho)I&C^{\top}\\ C&0\end{bmatrix}\begin{bmatrix}w\\ \eta\end{bmatrix}=\begin{bmatrix}A^{\top}b+\rho(z-u)\\ d\end{bmatrix}$,
where $\rho>0$ is the ADMM penalty parameter and $\eta$ is a vector of Lagrange multipliers for the equality constraints. The $z$-update is just componentwise clipping, $z^{k+1}=\Pi_{[-M,M]^{2R+1}}(w^{k+1}+u^{k})$, and the dual update is $u^{k+1}=u^{k}+(w^{k+1}-z^{k+1})$. We again monitor the objective, equality residual, and primal difference $\|w^{k+1}-w^{k}\|$.
5.3 Interior-point reference
Finally, we use the CVXPY modeling framework with a cone solver (SCS in our experiments) to obtain a reference solution to (19). This interior-point-based solution serves as a baseline for the quality of the first-order methods. The corresponding routine solve_cvx sets up the quadratic objective and linear constraints, and returns the optimal stencil $w_{\mathrm{cvx}}$ when CVXPY is available.
6 Numerical experiments
All experiments were implemented in Python using NumPy for linear algebra and SciPy for FFTs. The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5. The interior-point reference solution is computed using CVXPY with the SCS solver. Runtimes are reported for a single-threaded run on an Apple M4 laptop with 16 GB RAM using Python 3.11. Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
We now present numerical experiments in the 1D Maxwell setting to evaluate the learned stencils and the performance of the different solvers. All experiments use the Crank–Nicolson (CN) time integrator with the learned spatial operator.
6.1 Experimental setup
Unless otherwise stated, we use the following default parameters:
  • Spatial domain $[0,L]$ with $L=1$ and $N=64$ cells, $\Delta x=L/N$.
  • Training data: $n_{\mathrm{sims}}=200$ random spectral samples with $m_{\max}=5$ Fourier modes.
  • Regularization parameter $\lambda=10^{-6}$ and box parameter $M=100$.
  • Stencil radii $R\in\{1,2,3,4\}$, corresponding to stencil length $2R+1$.
  • Time integration: Crank–Nicolson (CN) applied to the semi-discrete system with time step $\Delta t=0.5\Delta x$ and $N_{t}=300$ time steps.
The initial condition for time integration is chosen as a single Fourier mode,
$E(x,0)=\sin(2\pi x/L),\qquad H(x,0)=\cos(2\pi x/L),$
and we compute a reference solution using the exact central-difference stencil
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$ (22)
We then integrate the Maxwell system with each learned stencil using CN and measure the relative $L^{2}$ error in the electric field at final time:
$\mathrm{err}(w)=\frac{\|\mathbf{E}^{\mathrm{learned}}(T)-\mathbf{E}^{\mathrm{FD}}(T)\|_{2}}{\|\mathbf{E}^{\mathrm{FD}}(T)\|_{2}}.$ (23)
We also compute the energy-constraint residual
$r_{\mathrm{eq}}(w)=\|Cw-d\|_{2},$
which is zero for an exactly skew-symmetric stencil.
As an additional diagnostic, we track the discrete electromagnetic energy
$\mathcal{E}^{n}:=\frac{1}{2}\Delta x\sum_{i}\bigl[(E_{i}^{n})^{2}+(H_{i}^{n})^{2}\bigr]$
over time. For the exact stencil and CN time-stepping, the code confirms that $\mathcal{E}^{n}$ is constant to within floating-point roundoff, consistent with the analysis in Section 2.
6.2 Learned stencils
Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$. We observe that all methods recover a stencil very close to the exact one, with small differences due to the finite training set and regularization.
Method $w_{-1}$ $w_{0}$ $w_{+1}$
Exact FD 32.000000-32.000000 $0.000000$ $32.000000$
PG 32.906110-32.906110 $0.000000$ $32.906110$
NAG 32.906098-32.906098 $0.000000$ $32.906098$
ADMM 32.906114-32.906114 $2.64\!\times\!10^{-16}$ $32.906114$
CVX 32.906115-32.906115 $1.57\!\times\!10^{-14}$ $32.906115$
Table 1: Learned stencils for radius $R=1$ compared with the exact central-difference stencil.
In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution. The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space-time evolution of the electric field $E(x,t)$ for the learned stencils and the exact central-difference stencil.
Figure 2: Final-time electric field $E(x,T)$ for the exact central-difference stencil and the learned stencils for $R=1$ under Crank–Nicolson time-stepping. The curves are visually indistinguishable at the scale shown.
This experiment serves as a sanity check for the learning framework: for the smallest radius $R=1$, the energy-constrained optimization recovers a stencil that is essentially indistinguishable from the classical centered difference, and the corresponding Crank–Nicolson time integration reproduces the reference solution to within plotting accuracy. In other words, when the true operator is known and low order, the learned discretizations do not degrade either accuracy or energy behavior, while providing the same skew-adjoint structure that underpins exact discrete energy conservation. This $R=1$ case therefore provides a baseline for the higher-radius stencils and noisy-training experiments considered below.
6.3 Spatial convergence for the ADMM–learned stencil
We consider the 1D Maxwell system in normalized units,
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E,$ (24)
with $\varepsilon_{0}=\mu_{0}=1$ so that the wave speed is $c=1$. On the periodic domain $x\in[0,L]$ we choose the traveling-wave exact solution
$E_{\mathrm{ex}}(x,t)=\sin\!\left(\frac{2\pi}{L}(x-t)\right),\qquad H_{\mathrm{ex}}(x,t)=\cos\!\left(\frac{2\pi}{L}(x-t)\right),$ (25)
which satisfies (24) and is $L$-periodic in space. The initial data at $t=0$ are taken from (25), and the numerical solution is advanced to a final time $T=10$ using the Crank–Nicolson time integrator with a spatial derivative given by the ADMM–learned stencil of radius $R=1$.
Let $E^{N}_{i}$ denote the numerical approximation to $E_{\mathrm{ex}}(x_{i},T)$ at the grid points $x_{i}=i\,\Delta x$, $i=0,\dots,N_{x}-1$. We measure the relative discrete $L^{2}$ error in the electric field by
$\mathrm{err}_{E}(\Delta x)=\frac{\left\|E^{N}-E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}}{\left\|E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}},\qquad\|v\|_{\ell^{2}}^{2}=\Delta x\sum_{i=0}^{N_{x}-1}v_{i}^{2}.$ (26)
The timestep is chosen proportionally to the mesh width, $\Delta t=0.2\,\Delta x$, so that we probe the spatial convergence of the learned stencil.
Table 2: Convergence of the ADMM–learned stencil with radius $R=1$. Relative $L^{2}$ error in $E(\cdot,T)$
$N_{x}$ $\Delta x$ rel. $L^{2}(E)$ error order
64 $1.562\times 10^{-2}$ $2.097\times 10^{0}$
128 $7.812\times 10^{-3}$ $4.875\times 10^{-1}$ $2.10$
256 $3.906\times 10^{-3}$ $1.192\times 10^{-1}$ $2.03$
512 $1.953\times 10^{-3}$ $585\times 10^{-2} 2.2$02.20
The table2 confirms approximately second-order spatial accuracy for the ADMM–learned radius-$1$ stencil.
Figure 3: Log–log plot of the relative $L^{2}$ error $\mathrm{err}_{E}(\Delta x)$ in $E(\cdot,T)$ versus $\Delta x$ for the ADMM–learned stencil.
6.4 Discrete energy conservation for learned Maxwell stencils
We next examine how closely the learned stencils preserve the semi-discrete electromagnetic energy under Crank–Nicolson time integration.
Figure 4: Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time stepping with the exact central-difference stencil and the structure-preserving learned stencils, shown on a linear (left) and logarithmic (right) scale.
Figure 4 compares the discrete electromagnetic energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time integration using the exact central-difference stencil and the structure-preserving learned stencils. On the linear scale, all curves are essentially flat and remain at the level of $10^{-12}$, indicating that the total energy is conserved up to machine precision for all methods. The logarithmic view reveals small differences: the PG, ADMM, and CVX stencils produce energy drift of the same order as the exact FD stencil, while NAG exhibits a slightly larger but still very small drift. These results confirm that enforcing the skew-symmetry constraints during training yields learned stencils that are effectively energy-preserving under Crank–Nicolson time stepping.
6.5 Dispersion analysis
To complement the end-time $L^{2}$ error metrics, we carry out a Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping. For a given stencil $w$ with Fourier symbol $\lambda(\theta)$, the one-step Crank–Nicolson update for each Fourier mode has eigenvalue
$\mu(\theta)=\frac{1+\tfrac{\Delta t}{2}\,\lambda(\theta)}{1-\tfrac{\Delta t}{2}\,\lambda(\theta)}.$
(a) Amplification error $|\mu(\theta)|-1$.
(b) Phase dispersion $\arg\mu(\theta)/\theta$.
Figure 5: Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping: (a) amplification error and (b) phase dispersion for learned vs. centered-difference stencils.
We plot the amplification factor $|\mu(\theta)|$ and the normalized phase $\arg\mu(\theta)/\theta$ as functions of $\theta\in(0,\pi]$, and compare the learned energy-constrained stencils with standard centered-difference stencils of the same radius $R$; see Figs. 5(a)5(b).
Figure 5(a) shows that $|\mu(\theta)|\approx 1$ for all learned stencils and all wavenumbers, essentially indistinguishable from the centered-difference curves and the continuous reference. This confirms that the fully discrete schemes are non-dissipative and preserve the discrete electromagnetic energy in practice, consistent with the skew-adjointness constraints imposed on $D$.
The phase plots in Fig. 5(b) provide a more stringent test of accuracy. For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference. Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties; see, e.g., [25, 2].
6.6 Learning a nonstandard skew-adjoint operator
To mimic an effective spatial operator arising from unresolved heterogeneity or subgrid physics, we consider a “nonstandard” convolution stencil $w^{\star}\in\mathbb{R}^{2R+1}$ with radius $R=2$ that is skew-adjoint (i.e., satisfies the constraints (11)) but does not coincide with the usual fourth-order centered difference coefficients. In this setting the analytic form of the differential operator is not available, but we can still generate high-fidelity training trajectories by evolving Maxwell’s equations using $w^{\star}$.
We then (i) learn an energy-constrained stencil $w_{\mathrm{QP}}$ by solving the quadratic program (20) with ADMM, and (ii) compare it to the standard centered-difference stencil $w_{\mathrm{CD}}$ of the same radius. Numerically we obtain
$\frac{\|w_{\mathrm{QP}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 6\times 10^{-8},\qquad\frac{\|w_{\mathrm{CD}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 2.5\times 10^{-1},$
so the learned operator recovers the nonstandard stencil to nearly machine precision, while the standard stencil incurs a $25\%$ relative coefficient error.
Figure 6(a) compares the final-time electric field $E(x,T)$ produced by $w^{\star}$, $w_{\mathrm{CD}}$, and $w_{\mathrm{QP}}$. The centered-difference stencil exhibits a systematic phase and amplitude error, whereas the learned stencil is indistinguishable from the nonstandard operator. At the same time, all three stencils remain essentially energy-preserving under Crank–Nicolson time stepping (Figure 6(b)), demonstrating that the proposed method can capture effective, nonstandard differential operators while retaining a discrete Poynting theorem.
(a) Final-time electric field $E(x,T)$ for nonstandard and learned stencils.
(b) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$.
Figure 6: Nonstandard vs. standard skew-adjoint operators: learned stencil $w_{\mathrm{QP}}$ compared with $w^{\star}$ and the centered difference $w_{\mathrm{CD}}$.
6.7 Noisy training and ill-posed least squares
To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (17). Specifically, we draw $\widetilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d. $N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$. We then compute three stencils: (i) the standard centered difference $w_{\mathrm{CD}}$, (ii) an unconstrained least-squares stencil $w_{\mathrm{LS}}$ obtained by minimizing (17) with no constraints, and (iii) an energy-constrained stencil $w_{\mathrm{QP}}$ obtained by solving the quadratic program (20) with the same noisy data.
Although $w_{\mathrm{LS}}$ achieves a very small training error, it violates the skew-adjointness constraints and develops very large high-frequency coefficients. When used in Crank–Nicolson time stepping, the resulting Maxwell semi-discretization is no longer energy-stable, and the discrete energy $\mathcal{E}^{n}$ blows up (Figure 7(a)). In contrast, the energy-constrained stencil $w_{\mathrm{QP}}$ remains exactly skew-adjoint by construction and yields stable dynamics with energy drift comparable to the hand-derived centered difference stencil.
Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators. The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the central-difference baseline, while the unconstrained LS stencil produces a completely corrupted field. This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
(a) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ under noisy training.
(b) Final-time electric field $E(x,T)$ for noisy training.
Figure 7: Noisy training and ill-posed least squares: comparison of unconstrained LS, central difference, and energy-constrained stencils.
Figure 8: Space-time plot of the electric field $E(x,t)$ under noisy training perturbations.
6.8 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 9 plots the objective value versus iteration and wall-clock time for each first-order method for a fixed radius. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
(a) Objective vs. iteration.
(b) Objective vs. runtime.
Figure 9: Convergence behavior of PG, NAG, and ADMM for a fixed stencil radius $R$. Left: objective value versus iteration count. Right: objective value versus wall-clock runtime.
Figure 9(a) shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ. PG exhibits the expected monotone, roughly geometric decrease. NAG achieves a much faster reduction in the first $\sim 20$ iterations, but with the characteristic oscillations of accelerated gradient methods before settling near the optimum. ADMM appears as an almost flat curve at the bottom of the plot because it solves a small KKT system at each iteration; in this setting, one or two ADMM iterations are enough to reach an objective value comparable to the final PG/NAG iterates.
Figure 9(b) plots the same objective values against wall-clock time. Because PG and NAG have very cheap iterations, their curves almost coincide and reach a low objective well below $10^{-3}$ in about $10^{-2}$ seconds for our problem size. ADMM has a higher per-iteration cost due to the linear solve, but it reaches essentially its final objective in a single step, so the total runtime is competitive with the first-order methods in this regime. For larger-scale problems, we expect the trade-off between cheap but many iterations (PG/NAG) and expensive but very few iterations (ADMM) to be even more pronounced.
7 Conclusion
We proposed an energy-conserving, data-driven discretization framework for the one-dimensional Maxwell system. By learning convolution stencils from spectral training data under a linear skew-symmetry constraint, we obtain discrete derivative operators that are guaranteed to be skew-adjoint and therefore exactly conserve semi-discrete electromagnetic energy. Numerical experiments show that the resulting schemes match the accuracy of standard central differences while preserving these structural properties. Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for learned stencils satisfying the skew-adjoint constraint.
8 Discussion and extensions
Our results show that simple linear constraints can encode key physics in data-driven discretizations: enforcing $Cw=d$ yields skew-adjoint spatial operators that exactly preserve a semi-discrete energy identity, while remaining close to standard centered differences in the homogeneous 1D setting.
To probe more realistic regimes, we also considered a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data). A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in higher dimensions, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
Acknowledgements
The author gratefully acknowledges Profs. Xiao Fu and Nathan Gibson for valuable guidance and feedback throughout this work.
Appendix: Spectral structure of the learned stencil
Theorem 3 (Spectral structure and modal energy conservation for the learned stencil).
Let $w^{\star}\in\mathbb{R}^{2R+1}$ be any solution of the quadratic program (19) (or the equality-constrained variant (20)), and let $D^{\star}$ be the associated periodic convolution operator defined by (2). Let $\mu^{\star}(\theta)$ denote the Fourier symbol of $D^{\star}$,
$\mu^{\star}(\theta):=\sum_{k=-R}^{R}w_{k}^{\star}\,\mathrm{e}^{\mathrm{i}k\theta},\qquad\theta\in[-\pi,\pi].$
Then the following hold:
  1. (i)
    Because $w^{\star}$ satisfies the linear constraint $Cw^{\star}=d$, it is skew-symmetric in the sense of (10), i.e.
    $w_{0}^{\star}=0,\qquad w_{-\ell}^{\star}=-w_{+\ell}^{\star},\quad\ell=1,\dots,R.$
    Consequently, the symbol $\mu^{\star}(\theta)$ is purely imaginary and odd:
    $\mu^{\star}(\theta)\in\mathrm{i}\mathbb{R},\qquad\mu^{\star}(-\theta)=-\mu^{\star}(\theta)\quad\text{for all }\theta\in[-\pi,\pi].$
  2. (ii)
    Consider the semi-discrete Maxwell system with the learned operator $D^{\star}$,
    $\frac{d}{dt}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix}=A^{\star}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix},\qquad A^{\star}=\begin{bmatrix}0&D^{\star}\\ (D^{\star})^{\top}&0\end{bmatrix}$,
    together with the discrete $L^{2}$ inner product (4) and energy (5). Then $A^{\star}$ is skew-adjoint and diagonalizable in the discrete Fourier basis. For each discrete wavenumber $\theta$, the corresponding $2\times 2$ block has eigenvalues
    $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|.$
    In particular, each Fourier mode evolves as a harmonic oscillator with frequency $|\mu^{\star}(\theta)|$.
  3. (iii)
    The discrete electromagnetic energy associated with the learned stencil can be written as a sum of modal contributions and is conserved mode by mode. If $\widehat{E}^{\star}(\theta,t)$ and $\widehat{H}^{\star}(\theta,t)$ denote the discrete Fourier coefficients of $\mathbf{E}(t)$ and $\mathbf{H}(t)$, then
    $\mathcal{E}(t)=\frac{\Delta x}{2}\sum_{\theta}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr),$
    and for each $\theta$ one has
    $\frac{d}{dt}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr)=0.$
    Thus the total discrete energy and each modal energy contribution are preserved exactly by the semi-discrete learned Maxwell system.
Proof.
Because $w^{\star}$ is a solution of (19) or (20), it satisfies the linear constraint $Cw^{\star}=d$ by construction. By Corollary 1, this is equivalent to the skew-symmetry conditions (10), which immediately imply the structure of $\mu^{\star}(\theta)$ and the identities in part (i) via the same calculation as in Theorem 2.
Skew-symmetry of $w^{\star}$ implies $(D^{\star})^{\top}=-D^{\star}$, so the block matrix $A^{\star}$ is skew-adjoint with respect to the discrete inner product (4). Since $D^{\star}$ is a periodic convolution, it is diagonalized by the discrete Fourier transform; the $2\times 2$ block for each mode $\theta$ has the form
$A^{\star}(\theta)=\begin{bmatrix}0&\mu^{\star}(\theta)\\ -\mu^{\star}(\theta)&0\end{bmatrix}$,
whose eigenvalues are $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|$, yielding part (ii).
Finally, unitarity of the discrete Fourier transform implies that the discrete $L^{2}$ norms of $\mathbf{E}$ and $\mathbf{H}$ are equal to the sums of squared modal amplitudes, so the energy $\mathcal{E}(t)$ decomposes as in part (iii). For each fixed $\theta$, the $2\times 2$ system governed by the skew-Hermitian matrix $A^{\star}(\theta)$ preserves the Euclidean norm of $(\widehat{H}^{\star}(\theta,t),\widehat{E}^{\star}(\theta,t))$, which implies modal energy conservation and hence conservation of the total discrete energy. ∎